OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of each rational function. f(x) = (x^2 + 10x + 24) / (x + 6)

OpenStudy (johnweldon1993):

Can you factor the top?

OpenStudy (johnweldon1993):

Here I'll get you started \[\large x^2 + 10x + 24 \space\text{becomes} \space (x + 4)(x + 6)\] Now we have \[\large \frac{ (x + 4)(x + 6) }{ (x + 6) }\] or when simplified... \[\large (x + 4)\] So what is our 0? \[\large x + 4 = 0\] what is 'x'...?

OpenStudy (johnweldon1993):

That will be the zero of the graph...to find the discontinuity's you find what values of 'x' make your denominator = 0 \[\large x + 6 = 0\] what is 'x' here?